OpenStudy (anonymous):

18. What is the solution of the equation? sqrt 2x+13 -5=x Can Someone HELP ME PLEASE!!!

OpenStudy (anonymous):

Are you gonna help me?

OpenStudy (mathmale):

Lauren: I need some clarification first: Which of the following was what you meant? \[\sqrt{2x}+13-5=x\] or \[\sqrt{2x+13}-5=x\] These two expressions are dramatically different. Use parentheses here to show what's under the radical sign and what's not: Sqrt(2x+13) - 5 = x.

OpenStudy (anonymous):

The second one

OpenStudy (mathmale):

Good. Thought so. Our next step is to eliminate the radical sign. You'll find this job to be a lot easier if you'd first add 5 to both sides of the equation. Would you do that, please?

OpenStudy (anonymous):

I need help wit this whole thing

OpenStudy (anonymous):

I really don't know what im doing when it comes to math

OpenStudy (anonymous):

But only with numbers 6,7,8,9,11,12,18 and 22

OpenStudy (mathmale):

Lauren, with some guidance I'm sure you'll do fine and that you'll build up confidence. I'd suggest you not make statements such as "I really don't know what I'm doing when it comes to math," because it's too easy for you to believe something like that. Back to the question: If Sqrt(2x+13) -5 = x, please add 5 to both sides of the equation and simplify it.

OpenStudy (mathmale):

Doing this makes the problem solution a lot easier. For now, take my word for that.

OpenStudy (mathmale):

You should end up with \[\sqrt{2x+13}=x+5.\] Does this agree with your result?

OpenStudy (anonymous):

Not really

OpenStudy (mathmale):

You started out with \[\sqrt{2x+13}-5=x.\] Adding 5 to both sides should give you the following: \[\sqrt{2x+13}-5+5=x+5.\] Can you agree that this simplifies to \[\sqrt{2x+13)}=x+5?\]

OpenStudy (anonymous):

Yes

OpenStudy (mathmale):

Good! The next step is to get rid of that radical (the square root operator). Know how? If so, please describe what you'd do.

OpenStudy (mathmale):

Have you done similar problems in the past? If so, apply what you learned back then.